### v1

Abstract
We propose a novel visual place recognition approach, VOP, that efficiently addresses occlusions and complex scenes by shifting from traditional reliance on global image similarities and local features to image overlap prediction. The proposed method enables the identification of visible image sections without requiring expensive feature detection and matching. By focusing on obtaining patch-level embeddings by a Vision Transformer backbone and establishing patch-to-patch correspondences, our approach uses a voting mechanism to assess overlap scores for potential database images, thereby providing a nuanced image retrieval metric in challenging scenarios. VOP leads to more accurate relative pose estimation and localization results on the retrieved image pairs than state-of-the-art baselines on a number of large-scale, real-world datasets. The code is available at https://github.com/weitong8591/vop.
Keywords:
image retrieval image overlap relative camera pose vision transformer
1 Introduction
Image retrieval algorithms, particularly those concerning visual place recognition (VPR), play an essential role in computer vision and robotics [17, 15] problems. Place recognition is defined as the task of identifying the approximate location where a query image was taken, given a certain tolerance level [43, 77, 19, 2, 28, 85, 78, 24, 27, 36, 6, 31, 83]. This capability is crucial for estimating the state of robots and is widely utilized in autonomous, unmanned aerial, terrestrial, and underwater vehicles, as well as recent AR/VR devices.
Typically, VPR is approached as an image retrieval problem, where the image to be localized is compared against a large database of posed images and, optionally, a 3D reconstruction of the scene. The most similar images retrieved from the database are used to estimate the precise location of the input query image, e.g., via local feature detection and matching. The complexity of VPR stems from a variety of factors such as perceptual aliasing, seasonal variations, changes in viewpoint, the presence of dynamic objects, illumination differences, occlusions, weather conditions, and intricate environments, as highlighted in the studies by Hong et al.[30], Doan et al.[20], and Subramaniam et al[72].
Recent approaches conceptualize VPR as a task of estimating image similarity, conducting retrieval through learned image embeddings generated by a feature extraction architecture augmented with an aggregation or pooling mechanism, the most notable method being NetVLAD [2]. In this framework, images from the database are selected for retrieval based on their similarity to the query image, for example, by optimizing the cosine similarity between their embeddings. One notable drawback of this method is its failure to address partial overlaps resulting from occlusions. In practice, occlusions can significantly reduce image similarity scores, despite the unobstructed portions of the scene potentially offering valuable cues for localization.
To be robust to occlusions and to provide a more nuanced understanding of the query image in relation to the database, we propose a novel approach for VPR. Instead of relying on global similarities or local features, we frame the problem as an overlap prediction, enabling us to methodically understand which parts of a particular image are visible without having to perform feature detection and matching. We approach the problem by generating embeddings for individual image patches rather than the entire image, thereby facilitating the efficient establishment of patch-to-patch correspondences. These matches are subsequently utilized within a voting mechanism to calculate an overlap score for potential database images, providing a nuanced metric for image retrieval amidst complex visual environments.
Illustrative examples are provided in Fig. 1. In these examples, AnyLoc [33], a state-of-the-art retrieval solution, processes a query image alongside two database images: one sharing the scene with the query albeit with occlusion and another from a different environment. AnyLoc erroneously attributes higher similarity to the unrelated image. In contrast, our proposed approach successfully identifies the correct image and the areas likely to overlap.
Figure 1: An outdoor and indoor image retrieval example of cases when image similarity prediction by the state-of-the-art AnyLoc [33] methods fails, while the proposed overlap prediction selects the correct database image, as indicated by the retrieval scores visualized above/below the database images. The middle column shows the same scene as the query (left) with heavy occlusion. The right column depicts a different location.
The contributions of this paper are summarized as follows:
  1. 1.
    A novel approach, transitioning from global image similarity to overlap assessment which exhibits enhanced robustness to challenges such as occlusions. The proposed approach is built on patch-level analysis by learning patch embeddings suitable for visual localization. It implements a single radius search across embeddings of image patches in the database.
  2. 2.
    The introduction of a robust voting mechanism for image retrieval which demonstrates improvement over conventional similarity-based methods.
Moreover, we identify a limitation inherent in the standard metric, namely recall, for evaluating image retrieval accuracy as done in recent studies [5, 33]. Recall quantifies the success of a method in retrieving correct database images without considering any application-specific evaluation metrics. Consequently, we opt for the Area Under the Curve (AUC) metric, specifically focusing on relative pose accuracy, to provide a more nuanced and relevant performance assessment.
2 Related Work
Image retrieval methods for VPR can be categorized into two groups, one globally representing the whole image and the other locally for keypoints.
Localization the process of determining the position and orientation of an agent within a pre-built map, is pivotal across various domains such as autonomous driving [29], drones [40], and augmented reality [13, 45]. The differentiation in localization techniques arises from their scene representation methods – be it through explicit 3D models [25, 26, 32, 38, 42, 46, 60, 63, 64, 66, 73, 86, 62], sets of posed images [87, 88, 56, 7], or implicitly via neural network weights [35, 79, 81, 34, 8, 9, 4, 14, 11, 47] – and their approach to camera pose estimation, whether by 2D-3D [25, 26, 60, 62, 64, 73, 86] or 2D-2D [88, 7] matches, or through a composite of base poses [35, 81, 34, 65, 56, 47]. Enhanced robustness is often sought through sequential image analysis rather than single-frame methods, with [50] adapting sequence matching to graph search and [80] developing an online sequence matching technique for real-time localization. In practice, localization comprises two main steps: place recognition and precise localization stages. Here, we focus on the first step, finding potential locations of a query image.
Place recognition is often cast as an image retrieval problem [23, 2, 6, 28, 31, 37, 53, 54, 83, 5, 33] that consists of two phases. In the offline indexing phase, a reference map represented by an image database is created. In the online retrieval phase, a query image, taken during a later traverse of the environment, is coarsely localized by identifying its closest matches within the reference map.
For measuring closeness, traditional methods like Bag of Visual Words [16] (BoW) identify key local image patches and store them as a codebook. BoW is often used in conjunction with 128-dimensional SIFT [44] descriptors both for map creation and image retrieval. These high-dimensional features often require approximate nearest neighbor (NN) search or GPUs to accelerate the search process. Showcasing the unbroken popularity of such approaches, they are still widely employed in applications, e.g., in ORB-SLAM [48, 49, 12].
As demonstrated by Zheng et al.[89], using CNNs for outperforming hand-crafted features in various applications, namely through compact (fixed-length) representations, has gradually become the norm. Most recent methods utilize learned embeddings generated by a feature extraction backbone with an aggregation or pooling head, the most notable method being NetVLAD [2]. In recent works, transformer-based retrieval methods [21] have been introduced, incorporating metric learning and attention mechanisms. These methods utilize local and global features, or a fusion of the two, to identify similar images.
Global retrieval methods integrate or combine handcrafted or learning-based local descriptors, such as SIFT [44] through the bag-of-visual-words model [55, 70], or employ alternative algorithms [76]. There is a growing trend towards learning-based techniques that devise improved methods for feature aggregation. NetVLAD [2] introduces an end-to-end trainable image representation extractor, employing weak supervision through soft assignment to cluster centers. Generalized Mean Pooling (GeM) [57] improves by introducing a learnable pooling layer that generalizes both max and average pooling, with a CNN serving as the backbone. Additionally, AP-GeM [59] enhances ranking accuracy by minimizing a relaxed global mean Average Precision (mAP) over a set of images. A CNN-based solution, CosPlace [5], utilizes the same CNN backbone and GeM pooling but approaches the place retrieval issue as a classification problem. MixVPR [1] transitions feature maps into a compact representation space through cascade MLP blocks. More recently, AnyLoc [33] emerges as a versatile method suitable for various image retrieval contexts, exploring CNN-based or Vision Transformer (ViT)-based foundational models and aggregation techniques to construct a vocabulary for searching in a large database.
Image reranking. Visual place recognition techniques have increasingly embraced a two-phase approach that pairs global retrieval with subsequent reranking based on local features. MultiVLAD [3] is proposed to use the maximum similarity between the query VLAD descriptors and the query as the matching score. Similarly, Razavian. et al. [69] rely on the average L2 distance of each query sub-patch to the reference image. Patch-NetVLAD [28], for instance, innovates by integrating NetVLAD with region-level descriptors and multi-scale patch tokens for enhanced geometric verification. It effectively merges the strengths of local and global descriptors, offering high resilience to changes in condition and viewpoint. RRT [75] and TransVPR [82] further this development by incorporating image-level supervision and attention-based mechanisms, respectively, focusing on spatial relationships and feature relevance within images. CVNet and $R^{2}$ormer [90] introduce methodologies for replacing traditional geometric verification with dense feature correlation and transformer-based reranking, focusing on the precise alignment of features and relevance of image pairs. SuperGlobal [68] simplifies the reranking process by relying on global features and K-nearest-neighbors aggregation, showcasing a move towards more scalable and computationally efficient VPR solutions. These advancements highlight a trend toward utilizing machine learning techniques, including deep learning and transformers, to refine the accuracy and efficiency of image retrieval.
3 Visual Overlap Prediction (VOP)
In this section, we describe the method proposed for efficiently predicting visual overlap in large image collections.
3.1 Problem Statement
In place recognition problems, the aim is to find images $I_{j}\in\mathbb{R}^{H\times W\times C}$ containing the same landmark or scene as the query $I_{i}\in\mathbb{R}^{H\times W\times C}$ and distinguishable from others. Usually, it is done by transforming the images to a high-dimensional embedding space and measuring the distances to select the top-$k$ most similar images. In this embedding space, an image $I_{i}$ is represented by a high-dimensional global $\mathbf{g}_{i}\in\mathbb{R}^{d_{g}}$ or local descriptors $\mathbf{d}_{i},\in\mathbb{R}^{n\times d_{l}}$, using encoder $f(\cdot)$, where $n\in\mathbb{N}$ denotes the number of local features, and $d_{g}\in\mathbb{N}$ is set to one for the whole image, $d_{l}\in\mathbb{N}$ to 256, 128, etc. The matching score, then, can be computed as the similarities of the descriptors, $s(I_{i},I_{j})=sim(\mathbf{d}_{i},\mathbf{d}_{j})$ or $sim(\mathbf{g}_{i},\mathbf{g}_{j})$, for example, as their cosine similarity.
Instead of relying on a single score that quantifies the similarity of two images, we focus on patch-level representations and formulate the matching score as the overlap between the images. Let us split image $I$ into a set $\mathcal{P}$ of rectangular patches by a uniform grid consisting of $n\times n$ rows and columns. Given sets of such rectangular patches $\mathcal{P}_{i}$ and $\mathcal{P}_{j}$ in the two images, we can define a patch-level overlap score as follows:
$o(p,q)=\text{volume}(\text{cone}(\mathbf{R}_{i},\mathbf{t}_{i},\mathbf{K}_{i},p),\text{cone}(\mathbf{R}_{j},\mathbf{t}_{j},\mathbf{K}_{j},q)),$ (1)
for each pair $(p,q)$, where $p\in\mathcal{P}_{i}$ and $q\in\mathcal{P}_{j}$ are the image patches, $\mathbf{R}\in\text{SO}(3)$ are the absolute rotations of the cameras, $\mathbf{t}\in\mathbb{R}^{3}$ are their absolute positions, $\mathbf{K}\in\mathbb{R}^{3\times 3}$ represents the intrinsic camera parameters, function $\text{cone}(\cdot)$ represents the 3D cone defined by the camera pose and the current patch, function $\text{volume}(\cdot,\cdot)$ measures the overlap of two posed cones in three dimensions. While this measure is easy to calculate given the camera poses and intrinsics, we observed that it does not reflect real-world applications, e.g., 3D reconstruction. Thus, we reformulate this measure as follows:
$o^{\prime}(p,q)=\sum_{X\in\mathcal{X}}\llbracket\text{inside}(\mathbf{K}_{i}(\mathbf{R}_{i}\mathbf{X}+\mathbf{t}_{i}),p)\rrbracket*\llbracket\text{inside}(\mathbf{K}_{j}(\mathbf{R}_{j}\mathbf{X}+\mathbf{t}_{j}),q)\rrbracket,$ (2)
[…]
The overlap of two entire images $I_{i}$ and $I_{j}$ is defined as the sum of overlap scores across match pairs as follows:
$O(I_{i},I_{j})=\sum_{p\in\mathcal{P}_{i}}\sum_{q\in\mathcal{P}_{j}}o^{\prime}(p,q)\approx\sum_{(p,q)\in\mathcal{C}}o^{\prime}(p,q),$ (3)
where set $\mathcal{C}=\{(p,q)\;|\;p\in\mathcal{P}_{i},q=\arg_{w}\max o^{\prime}(p,w)\in\mathcal{P}_{j},w\in\mathcal{P}_{j}\}$ is a set of patch-to-patch correspondences defined by assigning the most overlapping patch from the database image to each patch in the query image. Please note that while we could easily define image overlap directly by Eq. 2, such a formulation does not allow for the next steps that allow learning an overlap score.
However, this formulation still does not allow for retrieval, as in practice, the camera pose corresponding to the query image is unknown, and our aim is to estimate it. To do so, we further approximate the image overlap as follows:
$O(I_{i},I_{j})\approx\sum_{(p,q)\in\mathcal{C}}\rho(sim(p,q)),$ (4)
where $\rho:\mathbb{R}\to\mathbb{R}$ is a robust function. For example, $\rho$ can be defined as a voting mechanism as $\rho(x)=\llbracket x>\epsilon\rrbracket$, where $\epsilon$ is a manually set threshold and $x$ is the patch similarity $sim(p,q)$.
Given such a scheme to quantify image overlap via the similarity of a set of patch correspondences, the most similar image $I^{*}$ that we retrieve for query $I_{i}$ from a set of database images $\{I_{j}\},j\in 1,\dots,N$ is the one with the highest overlap, calculated as follows:
$I^{*}=\arg_{j\in 1,\dots,N}\max O(I_{i},I_{j}).$ (5)
In the scenario where multiple images need to be retrieved for a single query, we straightforwardly define a ranking by sorting images with high visual overlap. Thus, image retrieval from image collections becomes a patch-matching problem. Example matches are visualized in Fig. 2.
Figure 2: Patch matching examples in an image pair (column) with varying patch sizes (16, 64, and 256). When using 256 patches, the embeddings are directly provided by Dino V2 [51]. We used average pooling (16 and 64 patches) to calculate other resolutions. The numbers inside a particular patch denote the match index. The color overlay is calculated by applying PCA to the patch embeddings.
3.2 Learning Patch Matching via Patch Similarities
In order to define the similarity of image patches $p$ and $q$ in images $I_{i}$ and $I_{j}$, respectively, we aim to learn an embedding that allows for casting the problem as a descriptor matching. Thus, we aim to distill embeddings $e_{p}\in\mathbb{R}^{d}$ and $e_{q}\in\mathbb{R}^{d}$ such that $sim(p,q)=\delta(e_{p},e_{q})$ is high if and only if $p$ and $q$ are overlapping and zero otherwise. The function $\delta(\cdot)$ measures the similarity in the embedding space, e.g., as the cosine similarity.
Figure 3: The learning pipeline of the proposed Visual Overlap Prediction (VOP) method is shown, with a frozen ViT backbone [51], trainable encoder head, attention layers, and contrastive loss on the similarities. The input to the training is image pairs that are broken down into rectangular patches to learn patch-to-patch overlap scores.
Figure 4: The proposed Visual Overal Prediction (VOP) method at inference. Given an input query image and an image collection representing the reference map, the images are broken down into rectangular patches, the frozen backbone [51] extracts patch-level features, which are then fed into an encoder network, obtaining the final embeddings. The image retrieval is performed by running a radius search in the embedding space for each query patch to find potential database matches. The final overlap scores are determined by robust voting on the formed “query patch”-to-“database patch” correspondences. In practice, the map embeddings are obtained in an offline.
Pipeline for learning embeddings is visualized in Fig. 3. Given a pair of images $I_{i}$ and $I_{j}$, the problem is to learn embeddings $e_{p}$ and $e_{q}$ as defined previously. First, we obtain patch features by inputting the images into a Visual Transformer architecture. We use a frozen DINO V2 [51] backbone. These features are then fed into an encoder head, comprising one linear projection layer plus a few fully connected layers with GELU activation and dropout layer to encode the centers of the patches. Given the corresponding patch pair $(p,q)$ as a positive sample, we train this encoder to predict similar embeddings for $p$ and $q$. To do so, we employ self and cross-attention layers and contrastive learning on the similarities.
Contrastive Learning. We use contrastive learning to learn a joint embedding space for the image patches. To do so, we form patch pairs $(p,q)$. Real-world scenes are rarely static, e.g., objects move or undergo non-rigid deformations and illumination changes. To ensure that the learned embedding is robust to such temporal changes, we use image augmentation techniques as implemented in GlueFactory [41], such as adding random brightness, blur, flip, and noise.
For a query patch $p$, a set of candidate patches $\{q,q_{1},q_{2},...,q_{n}\}$ are provided for training, where $\{q_{1},q_{2},...,q_{n}\}$ act as $n$ negative samples, depicting patches that do not have a common field-of-view with the query patch. In our training pipeline, half of the negative samples come from images that overlap with the query image in general, but the selected patches do not. There are 30-90% negative patches in those generally overlapping image pairs. The remaining set of samples contains images capturing other scenes or other parts of the same scene.
We train our model by optimizing both a contrastive loss and an attention-based one as follows:
$\mathcal{L}=\alpha_{1}*\mathcal{L}_{\text{contrastive}}+\alpha_{2}*\mathcal{L}_{\text{attention}}.$ (6)
The contrastive loss is defined as follows:
$\mathcal{L}_{\text{contrastive}}=\frac{\sum_{i}^{n}\sum_{j}^{n}l_{\text{GT}}(\sigma-sim(e_{p},e_{q}))^{2}-\sum_{i}^{n}\sum_{j}^{n}(1-l_{\text{GT}})sim(e_{p},e_{q})^{2}}{n^{2}},$ (7)
where $l_{\text{GT}}$ is the ground-truth labels for each pair of patches, e.g., as the similarities between two patches from a image pair, $\llbracket sim(\mathbf{p}_{n_{i}},\mathbf{p}_{n_{j}})>\epsilon\rrbracket$, where $sim(\cdot)$ is the similarity measurement of the $d$-dimensional descriptors, i.e., cosine similarity with $\sigma=1$ as the margin. It penalizes the cases where there are high similarities on the patches but non-overlapping in the ground truth, i.e., no visible 3D points. We resize the input images to a fixed size. Therefore, the overall loss is averaged over $n^{2}$ tentative matching patches on one image pair.
Additionally, the self and cross attention layers $A(\cdot)$ are applied to select the relevant information and measure the similarity of two patch descriptors. Then we replace $sim(\cdot)$ by $A(\cdot)$ in Eq. 7, and the final loss, is computed by Eq. 6. The attentions are only used for training and do not change the descriptors in inference. Note that the same losses can be employed to project global descriptors, and the zero-overlapping image pairs can be used as the negative examples.
In order to provide means for quick prefiltering, we add a classification (CLS) token to the employed Transformer network that is trained to distill global image embeddings based on the image overlap scores, shown in Table 5. These tokens will be used to select a subset of potentially overlapping images given the query.
Constructing Training Data. Now, we introduce the visual supervision used to build GT labels for training. The ground-truth overlapping patches of the training image pairs are achieved following LightGlue [41] employing pixel-wise depth and relative pose. The ground-truth matching local patches are found in two steps: First, a set of dense pixel coordinates $c_{i}$ are detected in the first image $I_{i}$, and transformed to 3D points $X_{i}$. The obtained 3D point cloud is filtered by non-zero depths and transformed to the reference camera, i.e., $kp_{i2j}$, by the given transformation matrix. Given camera intrinsics, we project the 3D points $X_{i}$ to the camera planes and check for visibility. We make sure that only circle-consistent points are kept by using the projected point, calculating its 3D coordinates, and back-projecting to the original image to perform a visibility check again. The ground-truth matches are supposed to contain the points with a consistent depth [41]. This procedure provides the initial set of positive and negative pairs of matching image patches.
Next, given the visible pixel matches, we first check which patches they belong to and define the matching patches by having at least one visible pixel correspondence. We assume one-to-one matching relations, i.e., the patch with the most patches matches in the image pair will be chosen if there are multiple. Note that the visual supervision is built mutually from the image to the reference and vice versa.
3.3 Image Retrieval by Overlap Scores
Given a query image $I_{i}$ and a set of database images ${I_{j}}_{j=1,...,N}$, our goal is to efficiently determine the top-$k$ images for $I_{i}$ maximizing overlap. To do so, in the offline mapping phase of the method, we pre-generate the patch embeddings of all images in the database.
The retrieval pipeline is visualized in Fig. 4. In the online phase, we obtain the patch and, optionally, the global embeddings of the query image. Next, for each query image patch, we perform a radius search in the database of patch embeddings to determine all potentially overlapping candidates. Setting a radius threshold $\epsilon$ is easy as the employed cosine similarity falls between 0 and 1. We tune this threshold during training and use a fixed one in all our experiments. Note that k-nearest-neighbors algorithm is insufficient in our case as we would like to find all potentially overlapping patches to determine image overlaps. Such a radius search can be efficiently performed by standard bounding structures, e.g., kd-tree, and other implementation tricks. Optionally, we also use the global descriptor to prefilter the images to a shortlist.
To obtain the top-$k$ images, we perform voting such that the overlap calculation becomes
$O(I_{i},I_{j})\approx\sum_{(p,q)\in\mathcal{C}}\llbracket sim(p,q)>\epsilon\rrbracket,$ (8)
quantifying the image overlap as the count of overlapping patches inside the images, or the sum of the similarities of the descriptors as
$O(I_{i},I_{j})\approx\sum_{(p,q)\in\mathcal{C}}\max(sim(p,q)-\epsilon,0).$ (9)
Weighted Voting. To highlight the distinguishable query patches, we propose to weight the query patches from each image using a bag of visual words approach [16]. Specifically, we apply TF-IDF [58] weights to all the query patches, assigning lower weights to patches that frequently appear in queries but are rare in the matched database patches. As illustrated below, it is preferable to have more neighbor patches if fewer images share the same neighborhood with the query patch. For each patch, the weight $t_{i}$ is calculated as follows:
$t_{i}=\frac{n_{id}}{n_{d}}\log\frac{N}{n_{i}},$ (10)
where $N$ represents the total number of database images, $n_{d}$ is the total number of patches in each image, 256 in our case, with a patch size of (14, 14) on the resized image size of 224. Parameter $n_{id}$ is the number of neighbor patches in database images, defined as the cosine similarities of the patch descriptors within a specified radius, and $n_{i}$ denotes the number of database images that contain at least one neighbor of the query patch. Patches that frequently appear are deemed less important and thus are down-weighted. Weight $t_{i}$ for each query patch is incorporated into the overlap score computation as follows: $t_{i}*O(I_{i},I_{j})$. Note that the map weights are precomputed during the mapping phase.
4 Experimental Results
Our experiments concentrate on estimating two-view epipolar geometry for pairs of retrieved images. The model is trained on 153 scenes from the MegaDepth dataset [39], which includes Internet-sourced images and depths reconstructed via Structure from Motion (SfM) and multi-view stereo techniques. Beyond the MegaDepth testing scenes, we evaluate its generalization on 11 scenes from PhotoTourism [71], the 13 scenes from ETH3D [67], and on the InLoc dataset [74]. We benchmark the proposed method against several competitors, including the foundational DINO V2, CNN-based retrieval methods like NetVLAD [2] and CosPlace [5], as well as the state-of-the-art (SOTA) universal retrieval method AnyLoc [33] utilizing the same DINO V2 backbone as we do. Besides the global-only retrieval methods, we compare VOP with the SOTA reranking-based methods: ViT-based $R^{2}$ormer [90] and CNN-based Patch-NetVLAD [28]. Also, we show the results of CosPlace when fine-tuned on the same dataset we train on.
Training Details. The models underwent training on the 153 MegaDepth scenes, employing contrastive loss to differentiate between negative and positive patch pairs. Each scene contributes 150 positive and 150 negative image pairs to the training set, standardized to an image resolution of $224\times 224$ pixels. We use the DINO V2 pre-trained model with a patch size of $14$ and extract $256$-dimensional descriptors. Negative images, characterized by zero overlaps (e.g., sourced from distinct scenes), contrast with positive pairs, which are stratified into three groups based on overlap scores ranging from 0.1 to 0.7 for random selection. Notably, the contrastive loss framework is predicated on patch-level negative/positive distinctions rather than entire images. Following the methodology of LightGlue [41], patch-level labels are generated from depth information and visual features, utilizing the GlueFactory library [52] for implementation.
The training was conducted over 30 epochs with a batch size of 64, employing matching recalls as the primary criterion for best checkpoint selection and utilizing a learning rate of $41e-$ 4. The voting scheme and the radius for neighbor searching were optimized on the validation set, which included a single scene for this purpose. During inference, scenes “0015” (comprising approximately 0.4K images) and “0022” (encompassing approximately 0.5K images) from MegaDepth were designated for testing, enabling a focused evaluation of the model performance on these specific subsets.
[…]
Moreover, we report the average number of inliers across image pairs as detected by the retrieval methods. It is important to note that in scenarios where $k>1$, for each query, we follow the practical approach and select only the image pair that has the highest number of inliers. This comprehensive evaluation allows us to thoroughly assess the performance not just in retrieving relevant images but also in accurately estimating their relative poses.
4.1 Pose Estimation on Megadepth
Table 1 presents the results on the MegaDepth dataset, highlighting the performance of two variants of our method. For VOP, we select the model that maximizes the pose AUC on the validation set. For VOP-recall, we choose the model that maximizes retrieval recall. This experiment highlights that these are two different objectives, and “recall” should not be used for retrieval if the final goal is localization or reconstruction. Retrieval recall provides the most similar images, but they may not be suitable for pose estimation due to short baselines. The proposed method secures the highest scores for the @10 criterion and exhibits the lowest median pose errors, thereby suggesting that recall alone may not be a comprehensive metric for evaluating the efficacy of VPR methodologies. Notably, the VOP-recall variant demonstrates superior accuracy compared to other methods when tailored to enhance recall, which underscores the limitations of recall as a sole metric.
The underlined results mark the optimal configuration for relative pose estimation accuracy, evidencing the superior performance of the proposed VOP approach when utilizing the top-5 images. This configuration outperforms the global-only retrieval baselines (the first four rows) on this dataset, reinforcing the significance of selecting appropriate metrics and model configurations tailored to the specific demands of real-world downstream applications. Besides, rows 5 and 6 in Table 1 show our method outperforms the recent reranking approaches $R^{2}$ormer [90] and Patch-NetVLAD [28] on Megadepth. Both lead to comparable results in top-10 retrieval but worse in the others.
Recall (%) $\uparrow$ AUC@10 $\uparrow$ Med. pose error ($\downarrow$ # inliers $\uparrow$
Top $\rightarrow$ 1 5 10 1 5 10 1 5 10 1 5 10
AnyLoc [33] 16.7 24.4 29.5 60.8 65.3 64.1 2.38 2.26 2.35 220.5 273 287.5
CosPlace [5] 19.3 31.2 37.5 54.5 53.5 64.3 3.14 2.34 2.27 239.5 285.0 294.5
NetVLAD [2] 19.5 27.4 32.0 56.5 64.2 63.6 2.85 2.36 2.38 234.0 286.5 294.0
DINO V2 [51] 18.2 26.5 32.1 61.1 65.6 64.8 2.32 2.29 2.31 219.0 270.0 282.5
$R^{2}$ormer [90] 15.2 16.9 26.5 57.9 65.4 65.7 3.03 2.24 2.21 108.0 185.5 214.0
P-NetVLAD [28] 18.9 22.5 29.7 59.7 64.9 64.5 2.67 2.22 2.27 212.5 264.0 283.5
VOP 17.3 22.7 29.2 61.8 67.6 66.0 2.36 2.02 2.28 177.5 246.5 251.0
VOP-recall 22.4 31.1 37.6 59.0 64.0 64.1 2.56 2.42 2.49 222.9 286.4 296.2
Table 1: Relative pose estimation on the MegaDepth dataset on the top-$k$ retrieved images using different retrieval methods, with the best results in bold at each $k$ and the overall best underlined. Standard VOP is trained to maximize the AUC score of the relative pose errors. Version VOP-recall maximizes the recall during training.
AUC@10 $\uparrow$ med. pose error ($\downarrow$ #inliers $\uparrow$
Top $\rightarrow$ 1 5 10 1 5 10 1 5 10
AnyLoc [33] 88.9 90.2 90.2 0.76 0.78 0.77 373.2 425.1 438.0
CosPlace [5] 88.3 89.7 89.7 0.77 0.77 0.78 349.7 426.8 437.2
NetVLAD [2] 89.5 90.3 89.9 0.74 0.74 0.79 365.8 147.3 440.1
DINO V2 [51] 87.1 89.2 88.9 0.69 0.67 0.67 338.9 378.2 399.8
$R^{2}$ormer[90] 49.1 74.5 85.2 35.2 6.94 0.83 116.5 299.1 379.6
P-NetVLAD [28] 86.7 88.4 90.2 0.81 0.83 0.81 333.3 412.4 466.8
VOP 83.6 91.1 90.4 0.69 0.61 0.72 293.9 414.7 460.0
Table 2: Relative pose estimation on the ETH3D dataset [67] on the top-$k$ retrieved images using different retrieval methods, with the best results in bold at each $k$ and the overall best underlined. The proposed VOP is trained on the MegaDepth dataset.
4.2 Generalization Experiments
Moreover, we test the proposed method on multiple real-world scenes from the ETH3D [67] and PhotoTourism [71] datasets for relative pose estimation, and InLoc [74] for localization. We use the AUC-validated model from MegaDepth, i.e., no fine-tuning or retraining were done on the following datasets.
ETH3D. The ETH3D [67] benchmark consists of images with various viewpoints from indoor and outdoor scenes, with ground-truth poses reconstructed by laser point clouds. 13 scenes are used for testing, tens of images from each. All the images in the dataset are used as queries to select database images from the rest. Table 2 shows that the proposed VOP method is always among the top-performing methods, being either the best or second best across all configurations. The overall best results in each metric are always obtained by the proposed VOP. This highlights that VOP performs accurately even without retraining and, thus, is applicable for unseen scenes. Unfortunately, R2Former does not generalize well on ETH3D.
PhotoTourism. Table 3 reports the pose estimation results on 11 testing scenes from PhotoTurism [71], with 100 images in each. We outperform the SOTA competitors in terms of pose estimation accuracy, which shows that the overlap-retrieved images are beneficial in such a geometric pipeline estimating relative poses. Moreover, all Tables 123 show that the number of inliers is not an indicator of accuracy, as we result in fewer inliers while being more accurate than baselines. We have an exceptionally large performance improvement on the top-1 retrieval scores and on the top-5 score on the PhotoTourism dataset.
AUC@10 $\uparrow$ med. pose error ($\downarrow$ #inliers $\uparrow$
Top $\rightarrow$ 1 5 10 1 5 10 1 5 10
AnyLoc [33] 56.3 60.3 60.6 3.57 2.90 2.81 265.1 334.6 348.6
CosPlace [5] 53.4 60.8 61.2 3.78 2.91 2.87 291.5 347.0 354.1
NetVLAD [2] 55.0 59.3 60.3 3.75 3.10 2.94 283.9 340.0 348.1
DINO V2 [51] 61.1 63.1 63.0 3.17 2.82 2.75 266.0 338.6 356.3
$R^{2}$ormer [90] 60.6 67.8 67.9 4.49 2.36 2.35 180.0 241.9 277.6
P-NetVLAD [28] 61.9 62.4 61.8 2.76 2.65 2.76 261.9 315.2 339.4
VOP 62.5 67.5 65.4 2.63 2.00 2.32 224.5 286.1 281.3
Table 3: Relative pose estimation on the PhotoTourism dataset [71] on the top-$k$ retrieved images using different retrieval methods, with the best results in bold at each $k$ and the overall best underlined. VOP is trained on the MegaDepth dataset.
InLoc. Additionally, we ran all methods on the InLoc dataset [74] (without fine-tuning) using hloc [61] for localization to demonstrate their generalization capabilities in the case of a large domain gap (indoor/outdoor). To retrieve top-40 database images, we prefilter the top-100 by CLS tokens and rerank them by VOP. Columns ’InLoc-DUC1’ and ’InLoc-DUC2’ in Table 4 show the proposed VOP achieves high accuracy, showcasing its generalization abilities.
MegaDepth ETH3D PhotoTourism InLoc-DUC1 InLoc-DUC2 AVG
Metrics AUC@10 on top-5 $\uparrow$ recall@5 on top-40 $\uparrow$
AnyLoc [33] 65.3 90.2 60.3 74.7 75.6 73.2
Cosplace [5] 53.5 89.7 60.8 69.2 74.8 69.6
Cosplacee* [5] 65.8 79.7 65.9 41.4 29.0 56.4
NetVLAD [2] 64.2 90.3 59.3 65.7 71.0 70.1
DINO V2 [51] 65.6 89.2 63.1 63.6 71.0 70.5
VOP 67.6 91.1 67.5 72.2 77.1 75.1
Table 4: Relative pose estimation (AUC@10) on three benchmarks and indoor localization (recall@5) on InLoc. Cosplace* is fine-tuned on Megadepth dataset.
Fairness of Comparison. To demonstrate that the improvements from the proposed VOP do not stem from using different training data, we fine-tuned Cosplace [5] on MegaDepth. As shown in the fourth row (Cosplace*) of Table 4, fine-tuning improves the performance of Cosplace on relative pose estimation on Megadepth and PhotoTourism. However, it largely reduces its accuracy on the ETH3D and InLoc datasets.
4.3 Discussion on Patch-level Descriptors
DINO V2 Features. We investigated the number of patches to be used to split the images. The experiments were conducted on the testing scenes of MegaDepth, with all the images resized to $224$. From the DINO V2 (ViT-14) pretrained model with a patch size of $14^{2}$, we can extract 256 patch descriptors. Then, average pooling is applied to aggregate the patch descriptors to different numbers of patches, such as 4, 16, 64. When the patch number is set to 1, DINO V2 is applied directly to obtain a global image descriptor. Fig. 5 shows the global descriptors (#patches = 1) do not always perform well on the image retrieval problem, as indicated by having a lower recall than when 64 patches are used. This clearly demonstrates that patch-level features have the potential to improve visual localization. Similar to the experiments in the previous sections, the patch-level descriptors learned for our method enhance the original DINO V2 features.
Figure 5: Image retrieval evaluation by DINO V2 backbone with different numbers of patches, e.g., 256. The recalls, AUC@10 scores, and median pose errors in degrees are shown at top-1, 5, and 10 retrievals. When only one patch is used (orange bar), directly DINO V2 features are employed.
Validation Prefilter Augment Dropout AUC@10 $\uparrow$ Med. pose error ($\downarrow$ # inliers $\uparrow$
Top $\rightarrow$ 1 5 10 1 5 10 1 5 10
VOP Recall $\times$ $\checkmark$ $\times$ 57.4 66.0 63.7 2.87 2.17 2.34 232.5 262 285.5
$\checkmark$ $\checkmark$ $\times$ 57.5 65.2 64.5 2.96 2.19 2.31 235.5 272.5 288.5
AUC $\checkmark$ $\times$ $\times$ 61.9 65.1 64.7 2.37 2.19 2.30 209.0 272.5 282.0
$\times$ $\checkmark$ $\times$ 63.7 66.3 65.3 2.23 2.18 2.31 179.5 222.0 233.0
$\checkmark$ $\checkmark$ $\times$ 64.0 66.7 66.9 2.26 2.18 2.22 213.0 263.0 256.5
$\checkmark$ $\checkmark$ $\checkmark$ 61.8 67.6 66.0 2.36 2.02 2.28 177.5 246.5 251.0
Table 5: Relative pose estimation on the MegaDepth dataset [71] on the top-$k$ retrieved images using different configurations, with the best results in bold at each $k$ and the overall best underlined. The validation column refers to the VOP model used. Prefilter indicates if the CLS token was employed to shortlist the potential candidates before overlap prediction. Augment refers to whether data augmentation was used.
Ablations. Table 5 shows the relative pose estimation performance on the top-$k$ images retrieved by the model chosen from the best recall/AUC, with or w/o global prefilter (CLS tokens), and dropout layer. Testing sets of MegaDepth are used, as in Table 1. The CLS prefiltering improves the AUC scores. However, it increases the median pose errors marginally at the same time. As shown in the fifth row, data augmentation is essential in robustly learning the embeddings. Also, the last two rows in Table 4 and Table 5 show that the dropout layer improves the performance on Megadepth and helps with generalization for pose estimation on other data and indoor localization.
Training. To better understand the training process, Fig. 6 illustrates the training and validation losses on patch-level contrastive loss and the average similarity changing among different epochs on MegaDepth. It shows the contrastive loss helps to learn the embeddings of negative patches less similar and closer to positive ones, and it converges fast. The similarities shown are averaged over all positive/negative patches of the validation set indicated by the GT labels built from 3D reconstructions.
Figure 6: Contrastive losses on the training set (left) and validation (middle) over different epochs shown for negative or positive patch pairs. The right plot shows the average cosine similarities over different patch samples.
Run-time & Query Speed. As we reduce the dimensionality of the DINO v2 features to 256, the embeddings of all patches of each image need a total of 512 kB, while, for AnyLoc, the storage per image is 384 kB. While we require slightly more storage than AnyLoc [33], the difference is small. In addition, we compare the time of querying an image from the database of different sizes using AnyLoc or VOP in seconds. Prefiltering top 20 images by DINO V2 CLS tokens and running VOP for reranking to get top-1 out of 500 images cost 0.009 seconds, while 0.003 for AnyLoc. Querying top-1 from 5k images (prefiltered to 100), our method costs 0.03 seconds, while AnyLoc runs in 0.02s. Note that VOP needs much more time for radius search without prefiltering. We recommend using VOP as a reranking method in retrieval.
5 Conclusion
Our paper introduces a novel Visual Place Recognition (VPR) method focusing on patch-level overlap prediction to address challenging view changes and occlusions. Through extensive testing on the MegaDepth, PhotoTourism and ETH3D datasets, we demonstrate the effectiveness of the proposed VOP method. It outperforms state-of-the-art retrieval approaches in terms of relative pose accuracy. Critically, we demonstrate the inadequacy of recall as the sole metric for VPR performance, advocating for metrics aligned with real-world applications. By comparing configurations optimized for AUC score and recall, our results underline the necessity of metric selection tailored to specific VPR tasks. We believe this research marks a significant step forward in VPR, emphasizing the importance of patch-level analysis and metric consideration for improved accuracy and application relevance.